# RIPRAG: Hack a Black-box Retrieval-Augmented Generation Question-Answering System with Reinforcement Learning

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) systems based on Large Language Models (LLMs) have become a core technology for tasks such as question-answering (QA) and content generation. However, by injecting poisoned documents into the database of RAG systems, attackers can manipulate LLMs to generate text that aligns with their intended preferences. Existing research has primarily focused on white-box attacks against simplified RAG architectures. In this paper, we investigate a more complex and realistic scenario: the attacker lacks knowledge of the RAG system’s internal composition and implementation details, and the RAG system comprises components beyond a mere retriever. Specifically, we propose the RIPRAG attack framework, an end-to-end attack pipeline that treats the target RAG system as a black box, where the only information accessible to the attacker is whether the poisoning succeeds. Our method leverages Reinforcement Learning (RL) to optimize the generation model for poisoned documents, ensuring that the generated poisoned document aligns with the target RAG system’s preferences. Experimental results demonstrate that this method can effectively execute poisoning attacks against most complex RAG systems, achieving an attack success rate (ASR) improvement of up to 0.72 compared to baseline methods. This highlights prevalent deficiencies in current defensive methods and provides critical insights for LLM security research.
@@ 1 Introduction
LLMs have revolutionized the field of artificial intelligence, demonstrating remarkable capabilities in understanding and generating human-like text. However, their inherent limitation lies in the static nature of their parametric knowledge, which can become outdated or lack specificity for certain domains. RAG [1] has been proposed to mitigate this issue by equipping LLMs with access to an external, updatable database. This paradigm enhances the factuality and relevance of generated responses, particularly in critical applications like question-answering and content generation, by dynamically retrieving and grounding responses in pertinent information.
@@
Existing research on RAG poisoning attacks can be broadly categorized into white-box and black-box methods. White-box attacks [2, 3, 4, 5, 6] assume full knowledge of the RAG system’s architecture, and utilize gradient information to optimize poisoned texts for higher retrieval probability. However, their critical defect is the unrealistic assumption of a naive RAG pipeline (e.g., a single embedding model and LLM). Modern RAG systems [7] often employ complex retrieval strategies such as hybrid search or GraphRAG [8], where gradient information is inaccessible, rendering white-box methods ineffective. Conversely, current black-box attacks are often superficial. For example, PoisonedRAG [2] merely inserts the target query into the poisoned text, failing to leverage interactive feedback from the system. Topic-FlipRAG [9] relies on a surrogate open-source retriever, whose performance degrades significantly when it diverges from the target system’s actual retriever. Furthermore, these methods perform poorly in scenarios with a low poisoning rate, where the number of poisoned documents is significantly lower than the number of retrieved documents.
To overcome these limitations, we propose RIPRAG, a novel black-box attack framework that treats the target RAG system as an opaque oracle. Our key insight is to leverage RL to optimize the generation of poisoned documents by utilizing Interactive feedback from the black-box system, thereby achieving effective Poisoning. Specifically, RIPRAG interacts with the target system by injecting candidate documents and observing whether the attack is successful. This feedback, combined with a textual similarity reward, guides an RL agent to iteratively refine its poisoning strategy, effectively adapting to the unknown internal mechanics of the RAG system and maximizing the attack success rate even under challenging conditions.
@@
  • We propose RIPRAG, the first framework to apply Reinforcement Learning to the problem of attacking RAG systems. We have successfully addressed the issue of poor performance in the current RAG poisoning methods under low poisoning rate scenarios.
  • We introduce Reinforcement Learning from Black-box Feedback (RLBF), a training methodology that optimizes attack policies using only the success/failure signal from the target system.
@@
  • To the best of our knowledge, we are the first to evaluate attack methods on RAG systems equipped with advanced, targeted defense mechanisms, providing a more rigorous and realistic security assessment.
@@ 2 Related Works
2.1 Attacks on RAG System
Attack methods on RAG systems can be broadly divided into two categories, i.e., white-box attacking and black-box attacking. White-box attacking refers to methods that optimize their poisoned documents with the inner information of RAG systems, including using the gradient of the retriever [2, 4, 5, 6, 10] or using the score given by the retriever [3] to maximize the probability of being chosen by the retriever. However, in most cases, the inner part of RAG systems is not visible. Moreover, for more advanced retrieval methods like GraphRAG [8] or Modular RAG [7], computing the gradient of the retriever is impractical because it is not a simple neural network model. Current research on black-box approaches is limited. PoisonedRAG [2] enhances the similarity between the poisoned document and the target query by directly inserting the target query itself into the poisoned document. Topic-FlipRAG [9], on the other hand, leverages gradients from an open-source retriever to optimize the poisoned document, which is essentially still a white-box method. In contrast to prior work, our approach optimizes the poisoned document using only the outputs of the black-box system, making it a truly black-box attack method.
2.2 Defending on RAG System
To mitigate the risks of RAG systems being attacked, numerous defense methods have been proposed. These defense approaches can be broadly categorized into model-based methods and feature-based methods. Model-based methods detect poisoned tokens in retrieved texts by masking tokens that significantly contribute to the retriever’s output [11] or by clustering the activation information of the LLM [12]. Feature-based methods identify poisoned texts by analyzing word frequency density [13], applying elbow detection to retrieval [14], or clustering embedding vectors [15].
2.3 Reinforcement Learning from Human Feedback
RL has become a cornerstone technique for aligning LLMs with human preferences. The foundational approach, RLHF, was introduced by OpenAI in InstructGPT [16]. This framework has been widely adopted in models like GPT-4 [17], Qwen3 [18], and DeepSeek [19]. Subsequent research has expanded RLHF in several directions. Reinforcement Learning from AI Feedback (RLAIF) [20] reduces reliance on human annotators by using LLMs as preference labelers, demonstrating competitive performance in tasks like summarization and harmlessness.
To simplify the complex pipeline of RLHF, DPO [21] was proposed, which bypasses explicit reward modeling by deriving an optimal policy directly from preference data via a closed-form solution. Methods like SimPO [22] and RLOO [23] eliminate the need for a reference model, reducing memory overhead while maintaining performance. GRPO was initially used in Deepseek-Math [24] to help LLMs enhance their mathematical capabilities, but it was later widely applied in RLHF as well.
@@ 3.2 Attacker’s background knowledge
The database, retriever, and generator are three core components of an RAG system. State-of-the-art RAG systems often include additional components, such as rerankers and knowledge graphs. We assume that the attacker has no knowledge of the specific components within the RAG system, cannot access the parameters of any individual component, and is unaware of how these components are organized or interconnected. In other words, the attacker’s background knowledge is limited to only two facts:
@@ 4 Method
Figure 1: An overview of the proposed RIPRAG framework. RIPRAG adopts an end-to-end RL strategy: Firstly, the poisoning small language model (SLM) generates a group of poisoned documents. Then, the attacker injects poisoned documents into the database of the target RAG system. Finally, the attacker chats with the QA system to evaluate whether the attack is successful, using the result and BM25 between the question and poisoned documents to optimize the poisoning SLM with RLBF.
To effectively optimize adversarial text generation against black-box RAG systems, we propose RIPRAG, an end-to-end framework that enhances attack efficacy through a reinforcement learning mechanism with composite rewards. Figure 1 provides an overview of the proposed RIPRAG framework. This decoupling mitigates reward sparsity and policy instability, enabling stable convergence toward high-success-rate adversarial injections.
@@ 4.1 Reinforcement Learning from Black-box Feedback
Reinforcement Learning from Human Feedback (RLHF) has pioneered a methodology for model alignment by utilizing human evaluations as an opaque signal, optimizing policies through reinforcement learning without requiring access to the internal mechanisms of the evaluator. This approach demonstrates that effective learning can occur solely from external feedback, even when the source’s decision process is entirely unobservable. Inspired by RLHF, we introduce Reinforcement Learning from Black-box Feedback (RLBF), which adapts this core principle to adversarial contexts in black-box systems. RLBF harnesses the inherent opacity of external feedback channels to manipulate model behavior, treating any black-box system as a viable source for reinforcement signals.
The functional similarity between human evaluators and artificial black-box systems is fundamental to this adaptation. Humans provide feedback based on complex, subconscious cognitive processes that are inherently non-transparent, mirroring the behavior of engineered black boxes where internal operations are concealed by design. Both entities serve as evaluative interfaces that output scalar rewards or preferences without disclosing the rationale behind these outputs. Consequently, the reinforcement learning framework, originally developed for human feedback, naturally accommodates any system exhibiting such opaque characteristics.
@@ 4.2 Batch Relative Policy Optimization
Group Relative Policy Optimization (GRPO) is a widely utilized method in reinforcement learning. However, it faces critical inefficiencies when applied to RLBF-based adversarial scenarios. GRPO operates by generating multiple candidate responses per query and normalizing rewards exclusively within each response group to derive relative advantages. However, in the context of RAG attacks, the homogeneity induced by adversarial objectives renders intra-group relative advantages nearly imperceptible, resulting in vanishing gradients that severely compromise policy optimization.
To circumvent this limitation, we propose Batch Relative Policy Optimization (BRPO), which shifts the normalization scope from individual query groups to the entire batch of queries, thereby sustaining non-trivial gradient magnitudes and enabling effective adversarial learning under the RLBF framework. Following SimPO [22], we also omit the reference model used in GRPO. Formally, the BRPO loss function is defined as:
$\hat{A}_{i,j,t}$ $=\frac{\mathcal{R}^{(i,j)}_{\text{RL}}-mean(\mathcal{R}_{\text{RL}})}{std(\mathcal{R}_{\text{RL}})},$ (1)
$\tau_{i,j,t}$ $=\frac{\pi_{\theta}(D^{(i)}_{j,t}\mid q^{(i)},D^{(i)}_{j,<t})}{\pi_{\theta_{\text{old}}}(D^{(i)}_{j,t}\mid q^{(i)},D^{(i)}_{j,<t})}$, (2)
$\hat{\tau}_{i,j,t}$ $=\text{clip}(\frac{\pi_{\theta}(D^{(i)}_{j,t}\mid q^{(i)},D^{(i)}_{j,<t})}{\pi_{\theta_{\text{old}}}(D^{(i)}_{j,t}\mid q^{(i)},D^{(i)}_{j,<t})},1-\epsilon,1+\epsilon),$ (3)
$\mathcal{L}_{\text{BRPO}}$ $=-\frac{1}{|\mathbb{Q}|}\sum_{i=1}^{|\mathbb{Q}|}\frac{1}{M}\sum_{j=1}^{M}\frac{1}{|D^{(i)}_{j}|}\sum_{t=1}^{|D^{(i)}_{j}|}\min\left[\tau_{i,j,t}\hat{A}_{i,j,t},\hat{\tau}_{i,j,t}\hat{A}_{i,j,t}\right],$ (4)
@@ 4.3 Poisoning Reward Design
This section details the reward used in the RLBF process and how they are integrated into RLBF. Our goal is to optimize the poisoning SLM by leveraging feedback from the black box to improve the attack success rate of generated text. Following the design of PoisonedRAG, we designed two reward signals used in the RLBF process: the similarity reward $r^{(i,j)}_{\text{sim}}$ and the attack reward $r^{(i)}_{\text{suc}}$.
@@ 4.3.1 Similarity Reward
To address the challenge of sparse gradients in adversarial training when the primary attack reward becomes uninformative, we introduce a similarity reward as a dense intermediate signal. This reward comprises two orthogonal components, both leveraging the BM25 algorithm to compute textual similarity. For the $j$-th poisoned document of $i$-th query, the similarity reward $r^{(i,j)}_{\text{sim}}$ is defined as the weighted sum of two BM25-based metrics:
$r^{(i,j)}_{\text{sim}}=\min[\alpha,\text{BM25}(q^{(i)},D^{(i)}_{j}),\mathbb{I}(a^{(i)}_{\text{tgt}}\text{ in }D^{(i)}_{j})$], (5)
where $\alpha$ is the clipping coefficient to avoid reward hacking, $q^{(i)}$ is the target query, $D^{(i)}_{j}$ is the generated poisoned document, $a^{(i)}_{\text{tgt}}$ is the target answer, and $\mathbb{I}(\cdot)$ is the indicator function yielding $1$ when the target answer $a^{(i)}_{\text{tgt}}$ appears in document $D^{(i)}_{j}$ and 0 otherwise. The first term $\text{BM25}(q,D^{(i)}_{j})$ ensures semantic coherence with the user query, preventing degenerate outputs. The second term $\text{BM25}(a^{(i)}_{\text{tgt}},D^{(i)}_{j})$ steers generation toward lexical proximity with the target answer, directly facilitating attack success.
@@ 4.3.2 Attack Reward
$r^{(i)}_{\text{suc}}=\mathbb{I}\left(\mathcal{M}_{\text{RAG}}(q^{(i)},D^{(i)}_{1,\cdots,M})=a^{(i)}_{\text{tgt}}\right)$, (6)
where $\mathcal{M}_{\text{RAG}}$ denotes the black-box target RAG system, and $\mathbb{I}(\cdot)$ is the indicator function yielding $1$ upon successful attack execution and 0 otherwise. This binary formulation establishes a clear success criterion: the policy receives positive reward if and only if the generated adversarial documents $D^{(i)}_{1,\cdots,M})$ cause the target system to output the exact target response $a^{(i)}_{\text{tgt}}$.
@@
$\mathcal{R}_{i,j}=\lambda r^{(i)}_{\text{suc}}+(1-\lambda)r^{(i,j)}_{\text{sim}}$, (7)

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) systems based on Large Language Models (LLMs) have become a core technology for tasks such as question-answering (QA) and content generation. RAG poisoning is an attack method to induce LLMs to generate the attacker’s expected text by injecting poisoned documents into the database of RAG systems. Existing research can be broadly divided into two classes: white-box methods and black-box methods. White-box methods utilize gradient information to optimize poisoned documents, and black-box methods use a pre-trained LLM to generate them. However, existing white-box methods require knowledge of the RAG system’s internal composition and implementation details, whereas black-box methods are unable to utilize interactive information. In this work, we propose the RIPRAG attack framework, an end-to-end attack pipeline that treats the target RAG system as a black box and leverages our proposed Reinforcement Learning from Black-box Feedback (RLBF) method to optimize the generation model for poisoned documents. We designed two kinds of rewards: similarity reward and attack reward. Experimental results demonstrate that this method can effectively execute poisoning attacks against most complex RAG systems, achieving an attack success rate (ASR) improvement of up to 0.72 compared to baseline methods. This highlights prevalent deficiencies in current defensive methods and provides critical insights for LLM security research.
@@ 1 Introduction
RAG [1] has been proposed to mitigate the inherent limitation of LLMs, which lies in the static nature of their parametric knowledge that can become outdated or lack specificity for certain domains. By equipping LLMs with access to an external, updatable database, this paradigm enhances the factuality and relevance of generated responses, particularly in critical applications like question-answering and content generation, through dynamic retrieval and grounding of responses in pertinent information.
@@
Existing research on RAG poisoning attacks can be broadly categorized into white-box and black-box methods. White-box attacks [3, 4, 5, 6, 2] assume full knowledge of the RAG system’s architecture, and utilize gradient information to optimize poisoned texts for higher retrieval probability. However, their critical defect is the unrealistic assumption of a naive RAG pipeline (e.g., a single embedding model and LLM), which fails to account for modern RAG systems [7] that often employ complex retrieval strategies such as hybrid search or GraphRAG [8], where gradient information is inaccessible, thereby rendering white-box methods ineffective. Conversely, black-box methods do not rely on gradients of the target RAG system. Some of the methods [2] insert the target query into the poisoned text to improve retrieval probability, failing to leverage interactive feedback from the system. Others [9, 10] rely on a surrogate open-source retriever, while performance degrades significantly when it diverges from the target system’s actual retriever. Furthermore, these methods perform poorly in scenarios with a low poisoning rate, where the number of poisoned documents is significantly lower than the number of retrieved documents.
To overcome those limitations, we propose RIPRAG, a novel black-box attack framework that treats the target RAG system as an opaque oracle. Our key insight is to leverage RL to optimize the generation of poisoned documents by utilizing Interactive feedback from the black-box system, thereby achieving effective Poisoning. Specifically, RIPRAG interacts with the target system by injecting candidate documents and observing whether the attack is successful. This feedback, combined with a textual similarity reward, guides an RL agent to iteratively refine its poisoning strategy, effectively adapting to the unknown internal mechanics of the RAG system and maximizing the attack success rate even under challenging conditions.
@@
  • We propose RIPRAG, the first framework to apply Reinforcement Learning to the problem of attacking RAG systems. We use RL to enable an SLM to learn the interaction information of a black-box RAG system, thereby improving its performance under low poisoning rate scenarios.
  • We propose Reinforcement Learning from Black-box Feedback (RLBF), a novel RL paradigm that learns to optimize black-box systems using only input-output queries. Unlike standard RL settings, which assume access to environment internals or dense reward signals, RLBF operates under the more practical and challenging constraint of a completely opaque feedback mechanism.
@@
  • Most of the work on RAG security focuses on attacking vanilla or weakly defended systems. However, real-world deployments are increasingly protected. Our contribution lies in shifting the evaluation paradigm: To the best of our knowledge, we are the first to rigorously benchmark attack methods specifically against RAG systems equipped with advanced, targeted defenses. This provides a more realistic and practically relevant measure of their security posture.
@@ 2 Related Works
2.1 White-box Attacks on RAG System
White-box attacking refers to methods that optimize their poisoned documents with the inner information of RAG systems, including using the gradient of the retriever [2, 4, 5, 6, 11] or using the score given by the retriever [3] to maximize the probability of being chosen by the retriever.
However, in most cases, the inner part of RAG systems is not visible. Moreover, for more advanced retrieval methods like GraphRAG [8] or Modular RAG [7], computing the gradient of the retriever is impractical because it is not a simple neural network model.
2.2 Black-box Attacks on RAG System
Current research on black-box approaches is limited. PoisonedRAG [2] enhances the similarity between the poisoned document and the target query by directly inserting the target query itself into the poisoned document. Topic-FlipRAG [9], on the other hand, leverages gradients from an open-source retriever to optimize the poisoned document.
However, none of these methods effectively utilizes interaction information with black-box systems. The use of open-source retrievers is essentially an extension of white-box methods.
2.3 Reinforcement Learning
Reinforcement Learning (RL), with roots in optimal control and the Bellman equation, has evolved from early dynamic programming and temporal-difference methods to modern deep RL algorithms that have achieved superhuman performance in complex domains. Recently, RL has become a cornerstone technique for aligning LLMs with human preferences. RLHF was first introduced by OpenAI in InstructGPT [12], which has been widely adopted in models like GPT-4 [13], Qwen3 [14], and DeepSeek [15]. Subsequent research has expanded RLHF in several directions. RLAIF [16] reduces reliance on human annotators by using LLMs as preference labelers, demonstrating competitive performance in tasks like summarization and harmlessness.
To streamline RLHF’s complex pipeline, DPO [17] bypasses explicit reward modeling by directly deriving an optimal policy from preference data. Methods like SimPO [18] and RLOO [19] eliminate the need for a reference model, reducing memory overhead while maintaining performance. GRPO was initially used in Deepseek-Math [20] to help LLMs enhance their mathematical capabilities, but it was later widely applied in RLHF as well.
@@ 3.2 Attacker’s background knowledge
The database, retriever, and generator are three core components of an RAG system. Advanced RAG systems often include additional components, such as rerankers and knowledge graphs. We assume that the attacker has no knowledge of the specific components within the RAG system, cannot access the parameters of any individual component, and is unaware of how these components are organized or interconnected. In other words, the attacker’s background knowledge is limited to only two facts:
@@ 4 Method
Figure 1: A flowchart of the proposed RIPRAG framework.
To effectively optimize adversarial text generation against black-box RAG systems, we propose RIPRAG, an end-to-end framework that enhances attack efficacy through a reinforcement learning mechanism with composite rewards. Figure 1 presents the RIPRAG framework for adversarial RAG poisoning. Starting from a target question-answer pair, the Poisoning SLM generates a poisoned document that misattributes the answer and injects it into the RAG database. During query processing, the RAG system retrieves this document, producing an incorrect answer. The RLBF module then optimizes the Poisoning SLM via a feedback-driven loop, using attack rewards and similarity rewards to iteratively refine the poisoning strategy.
@@ 4.1 Reinforcement Learning from Black-box Feedback
In traditional RL, an agent learns an optimal policy $\pi_{\theta}$ through environment interactions $\mathcal{E}$, where actions $a_{t}\sim\pi_{\theta}(\cdot|s_{t})$ induce state transitions $s_{t}\rightarrow s_{t+1}$ and scalar rewards $r_{t}=r(s_{t},a_{t},s_{t+1})$ guide policy updates. RLBF redefines this paradigm for adversarial manipulation of black-box systems. The target system (e.g., commercial API, closed-source model) acts as an opaque environment $\mathcal{E}_{\text{bb}}$, while the adversary employs a generative policy $\pi_{\phi}$ to craft inputs $\mathbf{x}_{t}$ that steer $\mathcal{E}_{\text{bb}}$’s behavior. Rewards are derived implicitly from $\mathcal{E}_{\text{bb}}$’s feedback: outputs serve as reinforcement signals via a surrogate reward function $\hat{r}(\mathbf{y}_{t})$, where $\mathbf{y}_{t}=\mathcal{E}_{\text{bb}}(\mathbf{x}_{t})$. This leverages the system’s opacity as an optimization channel, enabling policy updates through $\nabla_{\phi}J(\phi)=\nabla_{\phi}\mathbb{E}_{\mathbf{x}_{t}\sim\pi_{\phi}}[\mathcal{L}(\hat{r}(\mathbf{y}_{t}))]$ without gradient access or architectural knowledge. Thus, RLBF preserves the RL framework while operating solely via external feedback, transforming black-box systems into reward models.
Our RAG poisoning method, RIPRAG, implements RLBF. For each query $q^{(i)}$, input $\mathbf{x}_{t}=D^{(i)}_{j}$ is sampled from $\pi_{\phi}=\text{SLM}(q^{(i)})$. The black-box RAG system $\mathcal{E}_{\text{bb}}=\mathcal{M}_{\text{RAG}}$ processes $q^{(i)}$ against its poisoned database to produce response $\mathbf{y}_{t}$. Despite the target RAG system is opacity, $\mathbf{y}_{t}$ yields reward $\hat{r}(\mathbf{y}_{t})=r^{(i)}_{\text{suc}}$. Policy $\phi$ (i.e., SLM parameters) is optimized via $\mathcal{L}_{\text{BRPO}}$.
@@ 4.2 Batch Relative Policy Optimization
GRPO is a widely utilized method in reinforcement learning, yet it faces critical inefficiencies in RLBF-based adversarial scenarios due to the homogeneity of candidate responses induced by adversarial objectives. This often results in imperceptible intra-group reward differences and vanishing gradients, which severely hinder policy optimization.
To address this issue, we propose Batch Relative Policy Optimization (BRPO), a novel approach that performs reward normalization across the entire batch of queries rather than within individual groups. This design sustains meaningful gradient magnitudes and enables stable adversarial learning under the RLBF framework. In addition, BRPO eliminates the need for a reference model, simplifying the optimization process while maintaining effectiveness. Formally, the BRPO loss function is defined as:
$\hat{A}_{i,j,t}$ $=\frac{\mathcal{R}^{(i,j)}_{\text{RL}}-mean(\mathcal{R}_{\text{RL}})}{std(\mathcal{R}_{\text{RL}})}$ (1)
$\tau_{i,j,t}$ $=\frac{\pi_{\theta}(D^{(i)}_{j,t}\mid q^{(i)},D^{(i)}_{j,<t})}{\pi_{\theta_{\text{old}}}(D^{(i)}_{j,t}\mid q^{(i)},D^{(i)}_{j,<t})}$ (2)
$\hat{\tau}_{i,j,t}$ $=\text{clip}(\frac{\pi_{\theta}(D^{(i)}_{j,t}\mid q^{(i)},D^{(i)}_{j,<t})}{\pi_{\theta_{\text{old}}}(D^{(i)}_{j,t}\mid q^{(i)},D^{(i)}_{j,<t})},1-\epsilon,1+\epsilon)$ (3)
$\mathcal{L}_{\text{BRPO}}$ $=-\sum_{i=1}^{|\mathbb{Q}|}\sum_{j=1}^{M}\sum_{t=1}^{|D^{(i)}_{j}|}\frac{\min\left[\tau_{i,j,t}\hat{A}_{i,j,t},\hat{\tau}_{i,j,t}\hat{A}_{i,j,t}\right]}{|\mathbb{Q}|\cdot M\cdot|D^{(i)}_{j}|}$ (4)
@@ 4.3 Poisoning Reward Design
This section details the rewards used in the RLBF process and how they are integrated into RLBF. Our goal is to optimize the poisoning SLM by leveraging feedback from the black box to improve the attack success rate of generated text. Following the design of PoisonedRAG, we designed two reward signals used in the RLBF process: the similarity reward $r^{(i,j)}_{\text{sim}}$ and the attack reward $r^{(i)}_{\text{suc}}$.
@@ 4.3.1 Similarity Reward
To address the challenge of sparse gradients in adversarial training when the primary attack reward becomes uninformative, we introduce a similarity reward as a dense intermediate signal. For the $j$-th poisoned document of $i$-th query, the similarity reward $r^{(i,j)}_{\text{sim}}$ is defined as:
$r^{(i,j)}_{\text{sim}}=\min[\alpha,\text{Sim}(q^{(i)},D^{(i)}_{j}),\mathbb{I}(a^{(i)}_{\text{tgt}}\text{ in }D^{(i)}_{j})$] (5)
[…]
@@ 4.3.2 Attack Reward
$r^{(i)}_{\text{suc}}=\mathbb{I}\left(\mathcal{M}_{\text{RAG}}(q^{(i)},D^{(i)}_{1,\cdots,M})=a^{(i)}_{\text{tgt}}\right)$ (6)
where $\mathcal{M}_{\text{RAG}}$ denotes the black-box target RAG system, and $\mathbb{I}(\cdot)$ is the indicator function yielding $1$ upon successful attack execution and 0 otherwise. This binary formulation establishes a clear success criterion: the policy receives positive reward if and only if the generated adversarial documents $D^{(i)}_{1,\cdots,M}$ cause the target system to output the exact target response $a^{(i)}_{\text{tgt}}$.
@@
$\mathcal{R}_{i,j}=\lambda r^{(i)}_{\text{suc}}+(1-\lambda)r^{(i,j)}_{\text{sim}}$ (7)